# OccluGaussian: Occlusion-Aware Gaussian Splatting for Large Scene Reconstruction and Rendering

### v1

@@ 2.2 Camera Clustering
In camera pose estimation of large-scale scenes, partitioning methods are commonly utilized to cluster cameras and segment scenes for parallel processing, which helps to reduce the computational complexity [20, 43, 26, 52, 37]. Some of them [37, 52, 5, 56, 7] use graph cut algorithms to aggregate cameras acoording to their co-visibilities, e.g., Out-of-Core-BA [37] and COLMAP [44] use the Metis graph partitioner [20]. Others [26, 43] use clustering algorithms based on the features of images and 3D key points. However, these methods often adopt only image or matching features, which are not sufficient enough to avoid occlusions within the segmented regions, as shown in Fig. 6. In this paper, we propose a scene division technique that exploits both camera position and co-visibility information, which is occlusion-aware and capable to align with the scene layout to conduct partitioning.
@@ 4.1 Experimental Setup
Implementation details. The models are optimized for 90,000 iterations with densification from iteration 1,500 to 45,000 at intervals of 100. We adopt the appearance modeling in VastGaussian [28] to fit appearance variations across captured images. For clustering the view graph, we set a fixed initial cluster number $K=10$. We iteratively refine the number of clusters until the camera count in each cluster falls within $[M_{c}-\sigma_{c}M_{c},M_{c}+\sigma_{c}M_{c}]$, where $\sigma_{c}=0.5$ and $M_{c}$ is the average camera count across all clusters.
@@ 4.2 Result Analysis
In Tab. 3, we validate our method on occlusion-free scenes using the Mill-19 [49] and UrbanScene3D [30] datasets. Our method maintains a clear advantage, highlighting the generality of our approach. These results shows the effectiveness of our scene division strategy, which enhances the correlation among training cameras within each region and improves reconstruction quality.
@@ 6 More Details of the OccluScene3D Dataset
There are three scenes in the OccluScene3D dataset: Gallery, Canteen and ClassBuilding. All the videos are recorded by a mobile phone with the wide-angle mode and landscape orientation at a frame rate of 60 Hz. We uniformly extract 3% of the frames, and use COLMAP [44] to estimate the camera intrinsic and extrinsic parameters. More details are shown in Tab. 8.
@@
Scene Area ($m^{2}$) #Video Duration ($min$) #Image
@@
Figure 10: Scene division results by OccluGaussian on the Mill-19 and UrbanScene3D datasets.
@@ 7.1 Additional Quantitative Analysis
Scene division strategy analysis. We present a comparison of scene division on the OccluScene3D dataset and the Zip-NeRF dataset among VastGaussian [28], CityGaussian [32], Hierarchical-GS [22], DOGS [6], and OccluGaussian in Fig. 7 and Fig. 8. The results demonstrate that the scene divisions produced by OccluGaussian better align with scene layouts. Consequently, OccluGaussian achieves superior reconstruction quality, which is proven by performing 3DGS optimization with the same hyperparameters across different scene division strategies, as illustrated in Fig. 9. We also present the division results of OccluGaussion on the Mill-19 dataset [49] and UrbanScene3D dataset [30] in Fig. 10, which demonstrate that our method also generalizes well in aerial capture scenes without occlusions.
@@ 7.2 Additional Ablations
Cluster number refinement. To obtain our final scene division, we iteratively refine the result from the graph clustering algorithm by splitting or ignoring clusters until the camera count in each cluster falls within [$M_{c}-\sigma_{c}M_{c},M_{c}+\sigma_{c}M_{c}]$, where $M_{c}$ is the average camera count across all clusters. Here we explore the influence with different $\sigma_{c}$ choices, as shown in Fig. 11 and Tab. 13. The division results remain occlusion-aware, and similar performances are obtained. Note that without this refinement step, comparable reconstruction quality can still be achieved; however, more GPU resources are required as more regions need to be reconstructed.

### v2

@@ 2.2 Camera Clustering
In camera pose estimation of large-scale scenes, partitioning methods are commonly utilized to cluster cameras and segment scenes for parallel processing, which helps to reduce the computational complexity [20, 43, 26, 52, 37]. Some of them [37, 52, 5, 56, 7] use graph cut algorithms to aggregate cameras acoording to their co-visibilities, e.g., Out-of-Core-BA [37] and COLMAP [44] use the Metis graph partitioner [20]. Others [26, 43] use clustering algorithms based on the features of images and 3D key points. However, these methods often adopt only image or matching features, which are not sufficient enough to avoid occlusions within the segmented regions, as shown in Fig. 6. In this paper, we propose a scene division technique that exploits both camera position and co-visibility information, which is occlusion-aware and capable of aligning with the scene layout to conduct partitioning.
@@ 4.1 Experimental Setup
Implementation details. The models are optimized for 90,000 iterations with densification from iteration 1,500 to 45,000 at intervals of 100. We adopt the appearance modeling in VastGaussian [28] to fit appearance variations across captured images. For clustering the view graph, we set a fixed initial cluster number $K=10$. We iteratively refine the number of clusters until the camera count in each cluster falls within $M_{c}-\sigma_{c}M_{c},M_{c}+\sigma_{c}M_{c}]$, where $\sigma_{c}=0.5$ and $M_{c}$ is the average camera count across all clusters.
@@ 4.2 Result Analysis
In Tab. 3, we validate our method on occlusion-free scenes using the Mill-19 [49] and UrbanScene3D [30] datasets. Our method maintains a clear advantage, highlighting the generality of our approach. These results show the effectiveness of our scene division strategy, which enhances the correlation among training cameras within each region and improves reconstruction quality.
@@ 5 Conclusion
6 Training details
All experiments were conducted on RTX 4090 GPUs, except for Hierarchical-GS on V100s due to higher memory demand. The reported time includes only 3DGS optimization, excluding preprocessing steps. Our division strategy takes an average of 9.28 minutes on the OccluScene3D dataset and can be greatly accelerated by porting it from CPU to GPU in future work. For fair comparison, we used consistent settings for all methods based on their public codes (except DOGS, which requires distributed training beyond our GPU capacitance) on the MILL-19 and URBANSCENE3D datasets: dividing each scene into 9 blocks, downsampling all images by 4, and optimizing each block for 60k steps with 200-step densification interval
@@ 7 More Details of the OccluScene3D Dataset
There are three scenes in the OccluScene3D dataset: Gallery, Canteen and ClassBuilding. All the videos are recorded by a mobile phone with the wide-angle mode and landscape orientation at a frame rate of 60 Hz. We use COLMAP [44] to estimate the camera intrinsic and extrinsic parameters. More details are shown in Tab. 8.
@@
Scene Area ($m^{2}$) #Video Duration (minmin) #Image
@@
Scene Garden Kitchen
Metrics PSNR SSIM LPIPS PSNR SSIM LPIPS
Vanilla 3DGS 27.41 0.868 0.103 30.31 0.922 0.12
Ours (w/o division) 27.32 0.879 0.068 30.78 0.922 0.065
Ours (two-block division) 27.37 0.879 0.067 30.73 0.923 0.064
Table 10: Quantitative comparison on the Mip-NeRF 360 dataset.
@@
Figure 10: Different division results on aerial datasets.
@@ 8.1 Additional Quantitative Analysis
Scene division strategy analysis. We present a comparison of scene division on the OccluScene3D dataset and the Zip-NeRF dataset among VastGaussian [28], CityGaussian [32], Hierarchical-GS [22], DOGS [6], and OccluGaussian in Fig. 7 and Fig. 8. The results demonstrate that the scene divisions produced by OccluGaussian better align with scene layouts. Consequently, OccluGaussian achieves superior reconstruction quality, which is proven by performing 3DGS optimization with the same hyperparameters across different scene division strategies, as illustrated in Fig. 9. For aerial datasets without severe occlusions, our method remains effective. Unlike previous methods using fixed layouts (e.g., 3×3 grid), our division approach is not strictly grid-based. Thanks to our view graph clustering, it can adapt to the distribution of cameras. This enables a more balanced number of cameras across blocks, which is crucial for higher performance, as also demonstrated by CityGaussian and DOGS. Such benefit is especially noticeable in elongated capturing scenarios (e.g., Building, Rubble), but less so in square scenes (e.g., Sci-Art), as illustrated in Fig. 10.
We further compare vanilla 3DGS with our method on two scenes from the Mip-NeRF 360 dataset. We evaluate our method with both no-division and two-block division settings. As these scenes are relatively small and occlusion-free, the divided blocks contain nearly all cameras, showing little difference compared to vanilla 3DGS (see Tab. 10).
@@
Figure 14: Ablation study of border set cameras on GALLERY and BERLIN. We also report the performance on the test sets.
Figure 15: Scene division on the OcclusionScene3D-E dataset.
Figure 16: Quantitative Evaluation on the OcclusionScene3D-E Dataset. The first line shows the ground truth, while the second line presents the rendering results of our OccluGaussian method.
@@ 8.2 Additional Ablations
Cluster number refinement. To obtain our final scene division, we iteratively refine the result from the graph clustering algorithm by splitting or ignoring clusters until the camera count in each cluster falls within $M_{c}-\sigma_{c}M_{c},M_{c}+\sigma_{c}M_{c}]$, where $M_{c}$ is the average camera count across all clusters. Here we explore the influence with different $\sigma_{c}$ choices, as shown in Fig. 11 and Tab. 14. The division results remain occlusion-aware, and similar performances are obtained. Note that without this refinement step, reconstruction quality remains comparable but demands more GPU resources for additional regions.
@@
Training camera selection.The ablation study of border set cameras is shown in Fig. 14. Without the border set, Gaussian primitives can become excessively large or elongated, leading to floaters or artifacts at region boundaries.
Scene Area ($m^{2}$) #Video Duration (minmin) #Image
Study Room 500 39 105 4752
Hall 800 33 97 6000
Meeting Room 1500 36 127 10000
ClassBuilding B 1000 27 63 7603
ClassBuilding C 1000 25 76 6435
Parking Lot 500 38 78 6060
Table 15: Statistics of the OccluScene3D-E dataset with six real scenes.
Scene PSNR SSIM LPIPS #GS (M) FPS
Study Room 22.52 0.84 0.190 2.6 314.73
Hall 22.70 0.79 0.212 1.7 341.76
Meeting Room 27.00 0.90 0.12 3.5 407.51
ClassBuilding B 22.24 0.85 0.158 1.8 595.38
ClassBuilding C 23.90 0.88 0.123 1.4 420.94
Parking Lot 22.46 0.86 0.146 1.6 314.73
Table 16: Quantitative results on the OccluScene3D-E dataset.
Initial $K$ Final $K$ PSNR SSIM LPIPS
1 1 25.18 0.882 0.145
5 5 25.35 0.895 0.114
7 7 25.59 0.902 0.103
10 7 25.81 0.903 0.099
15 10 25.33 0.899 0.101
Table 17: Different initial clustering numbers $K$ on the Gallery scene of OccluScene3D dataset.
9 Extended Dataset
To further advance research in occluded scene reconstruction, we introduce an extended version of the OccluScene3D dataset named OccluScene3D-E. It encompasses six diverse scenes: Study Room, Hall, Meeting Room, ClassBuilding B, ClassBuilding C, and Parking Lot, with detailed information in Tab. 15. As demonstrated in Fig. 15, applying OccluGaussian’s scene division method to OccluScene3D-E yields results that closely align with the underlying scene structures. Tab. 16 details the quantitative metrics. Ablation study on initial clustering number $K$ is shown in Tab. 17. The rendering outcomes for each scene, illustrated in Fig. 16, collectively confirm the robustness and effectiveness of our scene division strategy, as well as its capacity to produce high-fidelity visual results. Dataset page: https://occlugaussian.github.io/OccluScene3D/index.html.